OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4.

OpenStudy (dumbcow):

working backwards from synthetic division: assuming a polynomial of the form x^3 +x^2 +x + c -2 | 1 1 1 c -2 2 -6 ---------------- 1 -1 3 -4 c-6 = -4 c = 2 --> x^3 +x^2 +x +2

OpenStudy (dumbcow):

it says 3rd degree polynomial though

OpenStudy (anonymous):

x^3 + 2x^2 - 4

OpenStudy (anonymous):

\[x^3+x^2+x+2 \]

jhonyy9 (jhonyy9):

x+2 * x2 =x3 +2x2 so you need add to this -4 for you can gett it like remainder - so the answer will be x3 +2x2 -4

OpenStudy (anonymous):

so it would be (2x^2+3x=z^2)(x+2)=x^3+2x^2-4 ?

OpenStudy (dumbcow):

there are many possible correct answers: --> x^3 -x +2

OpenStudy (anonymous):

A Mathematica solution is attached.